Advertisements
Advertisements
Question
Three integers are chosen at random from the first 20 integers. The probability that their product is even is
Options
2/19
3/29
17/19
4/19
Advertisements
Solution
17/19
Number of ways in which we can choose three distinct integers from 20 integers = 20C3= 1140
We know that, if we take three odd numbers, there product will always be an odd number.
Out of 20 consecutive integers, 10 are even and 10 are odd integers.
Number of ways in which we can choose three distinct odd integers from 10 odd integers= 10C3 = 120
P(product is even) = 1 - P(product is odd)
= \[1 - \frac{120}{1140} = \frac{1140 - 120}{1140} = \frac{1020}{1140} = \frac{17}{19}\]
APPEARS IN
RELATED QUESTIONS
If `2/11` is the probability of an event, what is the probability of the event ‘not A’.
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)
In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?
The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?
A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine
- P(2)
- P(1 or 3)
- P(not 3)
In a simultaneous throw of a pair of dice, find the probability of getting:
8 as the sum
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of odd numbers
In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second
In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11
In a single throw of three dice, find the probability of getting a total of 17 or 18.
Three coins are tossed together. Find the probability of getting exactly two heads
Two dice are thrown. Find the odds in favour of getting the sum 5.
What are the odds in favour of getting a spade if the card drawn from a well-shuffled deck of cards? What are the odds in favour of getting a king?
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| 0.5 | 0.35 | ..... | 0.7 |
If A and B are two events associated with a random experiment such that P(A) = 0.3, P (B) = 0.4 and P (A ∪ B) = 0.5, find P (A ∩ B).
If A and B are two events associated with a random experiment such that
P(A) = 0.5, P(B) = 0.3 and P (A ∩ B) = 0.2, find P (A ∪ B).
There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C
One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.
In a single throw of two dice, find the probability that neither a doublet nor a total of 9 will appear.
100 students appeared for two examination, 60 passed the first, 50 passed the second and 30 passed both. Find the probability that a student selected at random has passed at least one examination.
If the probability of A to fail in an examination is \[\frac{1}{5}\] and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is
A box contains 10 good articles and 6 defective articles. One item is drawn at random. The probability that it is either good or has a defect, is
Out of 30 consecutive integers, 2 are chosen at random. The probability that their sum is odd, is
A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is
A box contains 10 good articles and 6 with defects. One item is drawn at random. The probability that it is either good or has a defect is
A box contains 6 nails and 10 nuts. Half of the nails and half of the nuts are rusted. If one item is chosen at random, the probability that it is rusted or is a nail is
In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy 10 tickets.
